OpenStudy (anonymous):

square root of 75 over 2 + 3 square root of 2 over 4 =?

OpenStudy (mathstudent55):

Is this the question? \( \dfrac{\sqrt{75}}{2} + \dfrac{3 \sqrt{2}}{4} = \)

OpenStudy (anonymous):

Yes.

OpenStudy (mathstudent55):

You need to add two fractions. What do two fractions that are to be added need to have?

OpenStudy (anonymous):

Uh, Idk?

OpenStudy (mathstudent55):

A common denominator.

OpenStudy (mathstudent55):

The denominators are 2 and 4. What is the LCD of 2 and 4?

OpenStudy (anonymous):

2?

OpenStudy (mathstudent55):

No. The LCD is the "least common denominator." That means the smallest number that you can divide by both numbers evenly (without remainder). The LCD of 2 and 4 is 4, since 4 is the smallest number divisible by both 2 and 4.

OpenStudy (mathstudent55):

The right fraction already has the LCD. Now we need the LCD in the left fraction.

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

\(\dfrac{\sqrt{75}}{2} + \dfrac{3 \sqrt{2}}{4} =\) \(=\dfrac{2}{2} \times \dfrac{\sqrt{75}}{2} + \dfrac{3 \sqrt{2}}{4} \) \(=\dfrac{2\sqrt{75}}{4} + \dfrac{3 \sqrt{2}}{4} \) Now that we have the same denominator, we can add the fractions.

OpenStudy (anonymous):

alright.

OpenStudy (mathstudent55):

Now we add the fractions.

OpenStudy (anonymous):

how?

OpenStudy (mathstudent55):

\(=\dfrac{2\sqrt{75}}{4} + \dfrac{3 \sqrt{2}}{4}\) \(= \dfrac{ 2\sqrt{75} + 3 \sqrt{2}}{4}\) Now we can do one final step. We can simplify \(\sqrt{75} \).

OpenStudy (mathstudent55):

\(= \dfrac{ 2\sqrt{25 \times 3} + 3 \sqrt{2}}{4}\) \(= \dfrac{ 2\sqrt{25} \times \sqrt {3} + 3 \sqrt{2}}{4}\) \(= \dfrac{ 2 \times 5\sqrt {3} + 3 \sqrt{2}}{4}\) \(= \dfrac{ 10\sqrt {3} + 3 \sqrt{2}}{4}\)

OpenStudy (anonymous):

How did you get that?

OpenStudy (mathstudent55):

I did it step-by-step. Just go over it a few times until you understand it.